40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Oct 26 19:10:59 2015, gautam, Update, Green Locking, AUX PDH loop characterization 
    Reply  Tue Dec 8 23:24:08 2015, gautam, Update, Green Locking, Y end laser (Lightwave) PZT calibration Ycalib_8Dec.pdf
       Reply  Sun Dec 13 21:55:28 2015, gautam, Update, Green Locking, Y end laser (Lightwave) PZT calibration IMG_5972.JPGServoY_TF_13Dec2015.pdfDatanCode.zipPSD_916Hz.pdf
          Reply  Mon Dec 14 16:27:11 2015, gautam, Update, Green Locking, Y-end AUX PDH noise breakdown YAUX_NB_Dec2015.pdfPDH_errSig_Calib.pdf
             Reply  Mon Jan 4 16:45:11 2016, gautam, Update, Green Locking, Y-end AUX PDH noise breakdown ErrSigBreakdown.pdfcontrolSigBreakdown.pdfYEnd_PDH_OLTF.pdf
          Reply  Mon Jan 4 16:09:54 2016, gautam, Update, Green Locking, Y end laser (Lightwave) PZT calibration choosingExcFreqs.pdflaserPZTcalib.pdf
Message ID: 11715     Entry time: Mon Oct 26 19:10:59 2015     Reply to this: 11865
Author: gautam 
Type: Update 
Category: Green Locking 
Subject: AUX PDH loop characterization 

I began my attempts to characterize the PDH loops at the X end today. My goal was to make the following measurements:

  • Dark noise and shot noise of the PD
  • Mixer noise
  • Servo electronics noise 

which I can then put into my simulink noise-budget scheme for the proposed IR beat setup.

I've made an Optickle model of a simple FP cavity and intend to match the measured PDH error signal from the X end to the simulated error signal to get the Hz/V calibration. I'll put the plots up for these shortly.

With regards to the other measurements, I was slowed down by remote data-acquisition from the SR785 - I've only managed to collect the analyzer noise floor data, and I plan to continue these measurements during the day tomorrow. 

ELOG V3.1.3-